OpenStudy (atlchic):

Write an equation of this line: Slope-5 & Y-Intercept--15

OpenStudy (anonymous):

y=-5x-15

OpenStudy (anonymous):

sorry. y=5x-15

OpenStudy (anonymous):

slope goes in front of the "x" and the y-int is put at the end.

OpenStudy (atlchic):

Write an equation for this line: (-2,3) and m= -1/4

OpenStudy (anonymous):

k, one sec

OpenStudy (atlchic):

sure

OpenStudy (anonymous):

any specific form? like point-slope or slope-intercept form.

OpenStudy (atlchic):

just an equation

OpenStudy (anonymous):

ok, i'll give it to you in a few different way.s

OpenStudy (anonymous):

Point- slope form:\[(y-3) = -1/4(x+2)\]

OpenStudy (anonymous):

slope-intercept form: \[y=-1/4x+5/2\]

OpenStudy (atlchic):

write equation of line: (-1,7), (8,-2)

OpenStudy (anonymous):

hehe. You going to be able to do these on your test?

OpenStudy (atlchic):

yea

OpenStudy (anonymous):

\[(y2-y1) = m(x2-x1)\] one second while I plug everything in.

OpenStudy (atlchic):

k

OpenStudy (anonymous):

\[(-2-7) = m(8+1)\]

OpenStudy (anonymous):

that's not done yet.

OpenStudy (atlchic):

write point-slope, (5,-2), m=3

OpenStudy (anonymous):

m = -8/9

OpenStudy (atlchic):

next question

OpenStudy (anonymous):

\[(y-7)= -8/9(x+1)\] \[y=(-8/9)x+(55/9)\]

OpenStudy (anonymous):

sorry, that is the final answer for the previous question

OpenStudy (atlchic):

next question

OpenStudy (anonymous):

y=3x-17

OpenStudy (atlchic):

write in standard form: y+3=2(x-4)

OpenStudy (anonymous):

y-2x=-11

OpenStudy (atlchic):

i thought standard form was ax+by=c

OpenStudy (anonymous):

-2x+y=-11

OpenStudy (atlchic):

write in slope-intercept form: y-1=-2(x+5)

OpenStudy (anonymous):

slope intercept form: y=mx+b

OpenStudy (atlchic):

yea can u put that in slope-intercept for me

OpenStudy (anonymous):

y=-2x-9

OpenStudy (atlchic):

solve 3y-4>-37

OpenStudy (anonymous):

y>-11